# Semantic Contextualization of Face Forgery: A New Definition, Dataset, and Detection Method

### v1

@@ IV Semantics-Oriented Face Forgery Detection
In this section, we describe a new SO-detection method, drawing inspiration from the proposed definition of face forgery.
@@ IV-A Probabilistic Formulation
where $f_{i}(\bm{x})$ can be regarded as the raw score (i.e., confidence) of the detector $\bm{f}$ on the $i$-th label. $\mathbb{I}[\cdot]$ is an indication function that excludes two illegal label relations. If the $j$-th node is manipulated (i.e., $y_{j}=1$), it necessitates that at least one of its parents, with indices in $\mathcal{P}_{j}$, is also manipulated, ensuring that $\sum_{i}y_{i}\neq 0$. Conversely, if the $i$-th node is manipulated, at least one of its children, with indices in $\mathcal{C}_{i}$, should also be manipulated, leading to $\sum_{j}y_{j}\neq 0$.
@@
Given that our label hierarchy is small, $Z(\bm{x})$ can be computed exhaustively via matrix multiplication.
@@ IV-B Bi-level Optimization
Assume a training minibatch $\mathcal{B}_{\mathrm{tr}}=\{\bm{x}^{(k)},\bm{y}^{(k)},\mathcal{I}^{(k)}\}_{k=1}^{K}$, where $\bm{y}^{(k)}=\{0,1\}^{N}$ is the complete ground truth vector and $\mathcal{I}^{(k)}\subseteq\{0,1,\ldots,N\}$ is the index set of observed labels. Our goal is to optimize a differentiable function $\bm{f}(\bm{x};\bm{\theta})$, parameterized by a vector $\bm{\theta}$, that outputs the raw scores of all $N$ labels, based on which the joint probability $p(\bm{y}|\bm{x};\bm{\theta})$ in Eq. (2) can be computed. A straightforward optimization objective is to minimize the negative log joint likelihood of the observed labels:
@@
where we marginalize the unobserved labels (i.e., the second summation in the second equation) in computing the joint likelihood [76]. Despite mathematical elegance, Eq. (IV-B) does not facilitate prioritization of the primary task, namely detecting whether a face image is real or fake, resulting in suboptimal performance (as will be clear in Sec. V-C). Thus, we opt for the multi-task learning setting, and minimize an alternative loss function that represents a linear weighted sum of the negative log marginal likelihoods of the observed labels:
@@
where $\bm{\lambda}=[\lambda_{0},\lambda_{1},\ldots,\lambda_{N-1}]^{\intercal}$ is the loss weight vector, trading off the $N$ tasks. To prioritize the primary task and to automate the loss weight adjustment, we resort to bi-level optimization [39], in particular the Auto-$\lambda$ algorithm [77]. At the upper level, we only minimize the loss of the primary task with respect to $\bm{\lambda}$ on the validation minibatch $\mathcal{B}_{\mathrm{val}}$. At the lower level, we minimize the overall loss defined in Eq. (IV-B) with respect to the model parameters $\bm{\theta}$ on the training minibatch $\mathcal{B}_{\mathrm{tr}}$. This leads to the following bi-level optimization problem:
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. During optimization, we sample training and validation data to be different minibatches in the same training dataset. Solving Problem (IV-B) necessitates second-order derivatives, which can be heavy on computing and memory usage. As suggested in [77], we employ the finite difference method as the efficient approximator.
TABLE II: Results of thirteen face forgery detectors under the testing Protocol-1 and Protocol-2 on the combined main and additional test sets of FFSC. All methods are trained on the training set of FF++ (or its respective augmented versions). “✗” indicates that no such evaluations can be properly done. The top two results are highlighted in bold
@@ V Experiments
In this section, we first employ the proposed FFSC dataset as the test set for evaluating current face forgery detectors predicated on various “intended” features. We next compare FFSC with current face forgery datasets [28, 40, 24] in fostering generalization when utilized as training sets. Last, we highlight the superiority of our SO-detection method over the traditional binary and multi-class classification-based methods.
@@ V-A FFSC as the Test Set
We introduce two new testing protocols with the goal of facilitating a fine-grained assessment of face forgery detectors, yielding valuable insights into their relative strengths and weaknesses.
  • Protocol-1: Generalization to novel manipulation methods for the same face attribute.
  • Protocol-2: Generalization to novel face attributes.
TABLE III: AUC results of various base models retrained on different datasets. Intra-dataset results are omitted, as denoted by “–”
@@
We examine thirteen face forgery detectors, including Rossler19 [28], CNND [78], F3-Net [58], FFD [79], Patch-Forensics [55], Face X-ray [20], MADD [21], FRDM [60], Lip-Forensics [71], RECCE [69], SBI [56], ICT [22], and CADDM [23]. Rossler19, CNND, and Patch-Forensics are commonly compared baseline models. F3-Net and FRDM expose face forgeries by high-frequency analysis. Face X-ray and SBI learn to detect blending boundaries. FFD and MADD employ attention mechanisms to extract and focus on the most relevant features. RECCE and CADDM, respectively, adopt face reconstruction and manipulation localization as auxiliary tasks. Notably, Lip-Forensics and ICT rely primarily on high-level lipreading and face identity features, respectively, rather than signal-level analysis.
We adopt the prediction accuracy (i.e., Acc (%)) and the area under the receiver operating characteristic curve (i.e., AUC (%)) as the evaluation metrics. The results are shown in Table II. Despite the superior intra-dataset performance achieved by nearly all detectors, they struggle to generalize to novel manipulations that alter the same face attribute. This provides a strong indication that existing detectors rely heavily on manipulation-specific features that are less generalizable. Two exceptions are Face X-ray [20] and SBI [56] that deliver satisfactory results under Protocol-1, largely due to their ability to detect blending boundaries commonly found in expression and identity manipulations. However, their effectiveness diminishes under Protocol-2, especially for the gender and pose attributes, where blending techniques are not involved during alteration. Engineered to identify local visual artifacts, CADDM [23] performs remarkably under Protocol-2, suggesting that current techniques for manipulating age, gender, and pose still have room for improvement in terms of visual fidelity. Correspondingly, CADDM achieves subpar performance for the expression attribute, which focuses on local semantic changes with fewer noticeable artifacts across the entire face.
The performance of the semantics-based detectors Lip-Forensics [71] and ICT [22] is not exceptional under either Protocol-1 or Protocol-2. This could be attributed to the fact that these detectors only exploit a single semantic face attribute, and do not capture the intrinsic interactions of multiple face semantics and their relationship with local face regions.
@@ V-B1 Training Dataset Comparison
We compare the proposed FFSC dataset with three established face forgery datasets: FF++ [28], DFDC [40], and ForgeryNet [24] for retraining various face forgery detectors as base models. It is important to note that each dataset may admit a different training strategy. Specifically, FF++ and DFDC treat face forgery detection as a standard binary classification task, while ForgeryNet approaches it as a multi-class classification task. The FFSC dataset is used to train SO-detectors. To assess the cross-dataset generalization, we include two more datasets: DF-1.0 [48] and Celeb-DF [46] as suggested in [71, 60, 63, 23, 70].
Six representative face forgery detectors are selected as base models: Rossler19 [28], CNND [78], MADD [21], FRDM [60], RECCE [69], and CADDM [23]. It is important to highlight that some detectors were initially trained with auxiliary tasks, such as image reconstruction [69] and manipulation localization [23]. To adhere to the original implementation of each detector when retraining on FFSC, we incorporate the auxiliary loss (if any) into the objective of the bi-level optimization problem, as denoted by Eq. (IV-B). Other training procedures and hyperparameter settings strictly follow the original publication. One exception is CADDM, where we omit the multi-scale face swap module because it depends on a reference image that is not accessible during retraining.
The AUC results are shown in Table III, from which we find that SO-detectors retrained on FFSC significantly surpass those retrained on the other three datasets. We believe such superior cross-dataset generalization arises because of the label hierarchy of FFSC, which facilitates SO-detection of face forgeries. This approach encourages the detectors to disregard features specific to manipulation methods and instead focus on more transferable features related to face attributes. Nevertheless, the detectors retrained using FFSC marginally fall short of those using ForgeryNet [24] on the DFDC dataset. This discrepancy may stem from the domain shift attributable to divergent data augmentation techniques and varying photography conditions between FFSC and DFDC. ForgeryNet addresses this gap by employing a broader spectrum of data augmentations, including several that coincide with those used in DFDC. Nonetheless, some augmentations (e.g., highly random brightness adjustment, excessive blurring, and extreme compression) tend to damage major face semantics and are thus not label-preserving.
TABLE VI: AUC results of different face forgery detection methods, in which base models are trained on the training set of FFSC and tested on the other four datasets. The prefixes “B-” and “M-” denote the binary and multi-class classification, respectively
@@ V-B2 Training Method Comparison
To single out the role of our SO-detection method, we compare it against the corresponding base model with the original training strategy by fixing the training set to FFSC.
Table IV shows the Protocol-1 results on the additional test set of FFSC (see the descriptions in Sec. III-B), which examines the generalization to novel manipulations that alter the same attributes in the training set. The primary observation is that our SO-detection method significantly improves the efficacy of all base models for almost every face attribute. RECCE [69] performs satisfactorily on its own, without leveraging the label hierarchy in FFSC. This can likely be credited to the integration of a face reconstruction auxiliary task within RECCE, which promotes extracting semantic face features, thus enhancing generalizability across the same face attributes. Frequency-based FRDM [60] and distortion-based CADDM [23] both experience noticeable boosts in Acc and AUC, suggesting that the features derived from the SO-detection are either superior to or at least provide a beneficial complement to the signal-level features used in these methods.
Table V presents the Protocol-2 results, where we train all base models on four out of five face attributes in the training set of FFSC and test them on the remaining face attribute in the test set of FFSC. Similarly, we see consistent improvements in performance under the more challenging Protocol-2. Even though the test face attribute is unobserved, our SO-detection method successfully captures the relationship between face attributes by modeling the joint probability distribution in Eq. (1), therefore enabling the SO-detectors to generalize across different face attributes.
Moreover, we demonstrate the advantages of our SO-detection method by comparing it to the standard binary and multi-class classification counterparts in Table VI, where we train all base models on the training set of FFSC and test them on FF++ [28], DFDC [40], DF-1.0 [48], and Celeb-DF [46]. It is evident that our proposed SO-detection method is consistently better across nearly all base models and test datasets. This indicates that exploiting label hierarchy can steer SO-detectors towards learning more generalizable and semantically rich features, enhancing their generalization capabilities. Interestingly, multi-class classification tends to hinder the generalizability of the detectors, compared to the binary classification baseline. This suggests that multi-class classification appears to promote dependence on specific manipulation cues, which act as “shortcuts” for detecting face forgeries and lead to potential overfitting.
@@ V-C Ablation Studies
Figure 6: Illustration of the SO-detector (Rossler19 [28] as the base model) in making predictions at the face attribute and region levels.
@@
We next contrast different optimization strategies for leveraging the label hierarchy: 1) minimizing the negative log joint likelihood in Eq. (IV-B), 2) training with the fixed equal weights in Eq. (IV-B), 3) training with the fixed and learned optimal weights (produced by the bi-level optimization), and 4) training with dynamic weighting average [80] without prioritizing the primary task. As presented in Table VIII, direct optimization of the joint likelihood of all observed labels results in a performance comparable to fixed equal weighting, with neither method giving prominence to the primary task. The application of dynamic weighting average appears ineffective in learning to emphasize the primary task, yielding only a marginal improvement over fixed equal weighting. Furthermore, our findings indicate that employing fixed optimal weights determined through bi-level optimization falls short of matching the performance of the default bi-level optimization procedure. This underscores the significance of dynamic weighting adjustment in promoting the primary task during optimization (refer to Fig. 5).
@@
Benefiting from the proposed SO-detection method, we can compute the marginal probability of each face attribute/region being manipulated, which endows the SO-detectors with some degree of interpretability. For example, in Fig. 6 (b), the SO-detector discerns the face as fake, and meanwhile indicates a likely alteration of expression, pinpointing the mouth and lips as manipulated regions.
@@ VI Conclusion and Discussion
In this work, we have given face forgery a new definition from the perspective of face semantics. Working with this definition, we constructed a new face forgery image dataset, FFSC, and proposed an SO-detection method. Extensive experiments have validated the value of the proposed FFSC dataset as training and test sets, and the superiority of our method in improving the generalizability of face forgery detectors.
It is important to note that the current definition of face forgery is imperfect. An excellent example is that reducing the frame rate in a speech video of House Speaker Nancy Pelosi makes her sound sluggish and slurred. Despite the obvious falseness of the video, it would still be classified as real under our definition as it does not involve any modifications of semantic face attributes. To address this issue, a possible solution could be adding a psychological_response node at the global face attribute level, connected to a behavior node at the local face region level. Additionally, it is natural to expand the label hierarchy in Fig. 2 by adding other attribute nodes (e.g., race and attractiveness).
Moreover, our current definition of face forgery is limited to face manipulation methods. It would be beneficial to expand this definition to include face creation methods, such as those utilizing GANs [3], diffusion models [2, 81] and computer graphics [82]. However, this expansion presents significant challenges. First, technology has advanced to the point where we can render an existing face presented in a real physical scene with such realism that it is indistinguishable from one captured by a digital camera (see Fig. 7). Second, there is evidence that generative models can memorize training data [83], and techniques exist to retrieve these memorized images (see Fig. 8). In both cases, determining the authenticity of a face image requires heightened awareness or insight to perceive its true nature. Taking a step back, a more pragmatic extension of our definition is to include digital manipulation methods for face images to images of natural scenes that do not necessarily contain faces.
@@
Figure 8: Memorization results of a diffusion model trained on a small-sized CelebA dataset [85], as presented in [83]. The top row depicts training face images downsampled to $80\times 80$, and the bottom row displays face images generated by the diffusion model, which closely resemble those from the training set.
Acknowledgement
The authors would like to thank Mr. Weiran Zhao for his kind help in data collection.

### v3

@@ IV Semantics-Oriented Face Forgery Detection
In this section, we describe a new SO-detection method, drawing inspiration from the proposed face forgery definition.
@@ IV-A Probabilistic Formulation
where $f_{i}(\bm{x})$ can be regarded as the raw score (i.e., confidence) of the detector $\bm{f}$ on the $i$-th label. $\mathbb{I}[\cdot]$ is an indication function that excludes two illegal label relations. If the $j$-th node is manipulated (i.e., $y_{j}=1$), it necessitates that at least one of its parents with indices in $\mathcal{P}_{j}$ is also manipulated, ensuring that $\sum_{i}y_{i}\neq 0$. Conversely, if the $i$-th node is manipulated, at least one of its children with indices in $\mathcal{C}_{i}$ should also be manipulated, leading to $\sum_{j}y_{j}\neq 0$.
@@
Given that our label hierarchy is not complex (see Fig. 2(a)), $Z(\bm{x})$ can be computed exhaustively via matrix multiplication.
@@ IV-B Bi-level Optimization
Assume a training minibatch $\mathcal{B}_{\mathrm{tr}}=\{\bm{x}^{(k)},\bm{y}^{(k)},\mathcal{I}^{(k)}\}_{k=1}^{K}$, where $\bm{y}^{(k)}=\{0,1\}^{N}$ is the complete ground truth vector and $\mathcal{I}^{(k)}\subseteq\{0,1,\ldots,N-1\}$ is the index set of observed labels. Our goal is to optimize a differentiable function $\bm{f}(\bm{x};\bm{\theta})$, parameterized by a vector $\bm{\theta}$, that outputs the raw scores of all $N$ labels, based on which the joint probability $p(\bm{y}|\bm{x};\bm{\theta})$ in Eq. (2) can be computed. A straightforward optimization objective is to minimize the negative log joint likelihood of the observed labels:
@@
where we marginalize the unobserved labels (i.e., the second summation in Eq. (IV-B)) in computing the joint likelihood [83]. However, Eq. (IV-B) does not facilitate prioritization of the primary task, i.e., detecting whether a face image is real or fake. Thus, we opt for the multi-task learning setting and minimize an alternative loss function that represents a linear weighted sum of the negative log marginal likelihoods of the observed labels:
@@
where $\bm{\lambda}=[\lambda_{0},\lambda_{1},\ldots,\lambda_{N-1}]^{\intercal}$ is the loss weight vector, trading off the $N$ tasks. To prioritize the primary task and to automate the loss weight adjustment, we resort to bi-level optimization [39], in particular, the Auto-$\lambda$ algorithm [84]. At the upper level, we only minimize the loss of the primary task with respect to $\bm{\lambda}$ on the validation minibatch $\mathcal{B}_{\mathrm{val}}$. At the lower level, we minimize the overall loss defined in Eq. (IV-B) with respect to the model parameters $\bm{\theta}$ on the training minibatch $\mathcal{B}_{\mathrm{tr}}$. This leads to
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. During optimization, we sample training and validation data as different minibatches in the same training dataset. Solving Problem (IV-B) necessitates calculating second-order derivatives, which can be computationally and memory intensive. To address this, we employ the finite difference method as an efficient approximator [84].
TABLE II: Results of thirteen face forgery detectors under the testing Protocol-1 and Protocol-2 on the combined main and additional test sets of FFSC. All methods are trained on the training set of FF++ (or its respective augmented versions). “✗” indicates that no such evaluation can be properly done. The top two results are highlighted in bold
@@ V Experiments
In this section, we first employ the proposed FFSC dataset as the test set for evaluating current face forgery detectors. We next compare FFSC with current face forgery datasets [28, 40, 24] in fostering generalization as training sets. Last, we highlight the superiority of our SO-detection method over the traditional binary and multi-class classification-based detectors.
@@ V-A FFSC as the Test Set
We introduce two new testing protocols with the goal of facilitating a fine-grained assessment of face forgery detectors, yielding valuable insights into their relative performance.
  • Protocol-1: Generalization to novel manipulation methods for the same face attribute in the training set.
  • Protocol-2: Generalization to novel face attributes absent from the training set.
TABLE III: AUC results of various base detectors retrained on different datasets. Intra-dataset results are omitted, as denoted by “–”
@@
We examine thirteen face forgery detectors, including Rossler19 [28], CNND [85], F3-Net [61], FFD [86], Patch-Forensics [58], Face X-ray [20], MADD [21], FRDM [63], Lip-Forensics [77], RECCE [74], SBI [59], ICT [22], and CADDM [23]. Rossler19, CNND, and Patch-Forensics are commonly compared baseline models. F3-Net and FRDM expose face forgery by high-frequency analysis. Face X-ray and SBI learn to detect blending boundaries. FFD and MADD employ attention mechanisms to extract the most relevant features. RECCE and CADDM, respectively, adopt face reconstruction and manipulation localization as auxiliary tasks. Lip-Forensics and ICT rely primarily on high-level lipreading and face identity features, respectively, rather than signal-level analysis.
We adopt the prediction accuracy (i.e., Acc (%)) and the area under the receiver operating characteristic curve (i.e., AUC (%)) as the evaluation metrics. The results are shown in Table II. Despite the superior intra-dataset performance achieved by nearly all detectors, they struggle to generalize to novel manipulations that alter the same face attribute. This provides a strong indication that existing detectors rely heavily on manipulation-specific features that are less generalizable. Two exceptions are Face X-ray [20] and SBI [59] that deliver satisfactory results under Protocol-1, due to their ability to detect blending boundaries commonly found in expression and identity manipulations. However, their effectiveness diminishes under Protocol-2, especially for the gender and pose attributes, where blending techniques are not involved during manipulation. Engineered to identify local visual artifacts, CADDM [23] performs remarkably under Protocol-2, suggesting that current techniques for manipulating age, gender, and pose still have room for improvements in terms of visual fidelity. Correspondingly, CADDM achieves subpar performance for the expression attribute, which focuses on local semantic changes with fewer noticeable artifacts.
The performance of the semantics-based detectors Lip-Forensics [77] and ICT [22] is not exceptional under either Protocol-1 or Protocol-2. This could be attributed to the fact that these detectors only exploit a single semantic face attribute, and do not capture the intrinsic interactions of multiple face semantics and their relationships with local face regions.
@@ V-B1 Training Dataset Comparison
We compare the proposed FFSC dataset with three established face forgery datasets: FF++ [28], DFDC [40], and ForgeryNet [24] for retraining various face forgery detectors as base models. It is important to note that each dataset may admit a different training strategy. Specifically, FF++ and DFDC treat face forgery detection as a standard binary classification task, while ForgeryNet approaches it as a multi-class classification task. And the FFSC dataset is used to train SO-detectors. To assess cross-dataset generalization, we include two more datasets: DF-1.0 [48] and Celeb-DF [46].
Six representative face forgery detectors are selected as base models: Rossler19 [28], CNND [85], MADD [21], FRDM [63], RECCE [74], and CADDM [23]. It is important to highlight that some detectors were initially trained with auxiliary tasks, such as image reconstruction [74] and manipulation localization [23]. To adhere to the original implementation when retraining on FFSC, we incorporate the auxiliary loss (if any) into the objective of the bi-level optimization problem in Eq. (IV-B). One exception is CADDM, where we omit the multi-scale face swap module because it depends on a reference image that is not accessible during retraining.
The AUC results are shown in Table III, from which we find that SO-detectors retrained on FFSC significantly surpass those retrained on the other three datasets. We believe such superior cross-dataset generalization arises because of the label hierarchy of FFSC, which facilitates the SO-detection of face forgery. This approach encourages the detectors to disregard features specific to manipulation methods and instead focus on more transferable features related to face attributes. Nevertheless, the detectors retrained using FFSC marginally fall short of those using ForgeryNet [24] on the DFDC dataset. This discrepancy may stem from the domain shift attributable to divergent data augmentation techniques and varying photography conditions between FFSC and DFDC. ForgeryNet addresses this gap by employing a broader spectrum of data augmentations, including several that coincide with those used in DFDC. Nonetheless, some augmentations (e.g., highly random brightness adjustment, excessive blurring, and extreme compression) tend to damage major face semantics and are thus not label-preserving.
TABLE VI: AUC results of different face forgery detection methods, in which base detectors are trained on the training set of FFSC. The prefixes “B-” and “M-” denote the binary and multi-class classification, respectively
@@ V-B2 Training Method Comparison
To single out the role of our SO-detection method, we compare it against the corresponding base model with the original training strategy on the FFSC training set.
Table IV shows the Protocol-1 results on the additional test set of FFSC (see the descriptions in Sec. III-B), which examines the generalization to novel manipulations that alter the same attributes. The primary observation is that our SO-detectors significantly improve the efficacy of all base models for almost all face attributes. RECCE [74] performs satisfactorily on its own, without leveraging the label hierarchy in FFSC. This can likely be credited to the integration of a face reconstruction auxiliary task, which promotes extracting semantic face features, thus enhancing generalization across the same face attributes. Frequency-based FRDM [63] and distortion-based CADDM [23] both experience noticeable boosts in Acc and AUC, suggesting that the features derived from the SO-detection are either superior to or at least provide a beneficial complement to the signal-level features used in the original methods.
Table V presents the Protocol-2 results, where we train all base models on four out of five face attributes in the training set of FFSC and test them on the remaining face attribute in the test set of FFSC. Similarly, we see consistent improvements in performance under the more challenging Protocol-2. Even though the test face attribute is unobserved, our SO-detectors successfully capture the relationships between face attributes by modeling the joint probability distribution, therefore enabling the SO-detectors to generalize across different face attributes.
Moreover, we demonstrate the advantages of our SO-detection method by comparing it to the standard binary and multi-class classification counterparts in Table VI, where we train all base models on the training set of FFSC and test them on FF++ [28], DFDC [40], DF-1.0 [48], and Celeb-DF [46]. It is evident that the proposed SO-detection method is consistently better across nearly all base models and test datasets. This indicates that exploiting label hierarchy can steer SO-detectors toward learning more generalizable features. Interestingly, multi-class classification tends to hinder detection generalizability, compared to the binary classification baseline. This suggests that multi-class classification appears to promote learning of manipulation-specific cues, which act as “shortcuts” for face forgery detection, leading to potential overfitting.
@@ V-C Ablation Studies
Figure 6: Illustration of our SO-detector (Rossler19 [28] as the base model) in making predictions at the face attribute and region levels.
@@
We next contrast different optimization strategies for leveraging the label hierarchy: 1) minimizing the negative log joint likelihood in Eq. (IV-B), 2) training with the fixed equal weights in Eq. (IV-B), 3) training with the fixed and learned optimal weights (supplied by the bi-level optimization), and 4) training with dynamic weighting average [87] without prioritizing the primary task. As presented in Table VIII, direct optimization of the joint likelihood of all observed labels results in a performance comparable to fixed equal weighting, with neither method giving prominence to the primary task. The application of dynamic weighting average appears ineffective in learning to prioritize the primary task, yielding only a marginal improvement over fixed equal weighting. Furthermore, our findings indicate that employing fixed optimal weights determined through bi-level optimization falls short of matching the performance of the default bi-level optimization. This underscores the significance of dynamic weighting adjustment in promoting the primary task during optimization (see Fig. 5).
@@
(a) Gender manipulation
(b) Identity manipulation
(c) Pose manipulation
Figure 7: Visualization of Grad-CAM activation maps. Each subfigure shows the forged face (left), base Rossler19 [28] heatmap (middle), and SO-Rossler19 heatmap (right).
V-D Further Analysis
Benefiting from the proposed SO-detection method, we can compute the marginal probability of each face attribute and region being manipulated, which endows the SO-detectors with some degree of interpretability. For example, in Fig. 6(b), our method discerns the face as fake, and meanwhile indicates a likely alteration of expression, pinpointing the mouth and lip as manipulated regions.
Additionally, we use Grad-CAM [88] to visualize the activation maps of the base Rossler19 [28] and SO-Rossler19 in Fig. 7. The results demonstrate that the SO-detector consistently targets meaningful semantic face regions, whereas the baseline detector frequently depends on “cheap shortcuts,” such as background artifacts for detection.
@@ VI Conclusion and Discussion
In this work, we have given face forgery a new definition from the perspective of face semantics. Based on this definition, we constructed a new face forgery image dataset, FFSC, and proposed a SO-detection method. Extensive experiments have validated the promise of the proposed FFSC dataset as training and test sets, and the superiority of our method in improving the generalizability of face forgery detectors.
It is important to note that the current definition of face forgery is imperfect. One counterexample is that reducing the frame rate in a speech video of House Speaker Nancy Pelosi makes her sound sluggish and slurred. Despite the obvious falseness of the video, it would still be classified as real under our definition as it does not involve any modifications of semantic face attributes. To address this issue, a possible solution could be adding a psychological_response node at the global face attribute level. Additionally, it is natural to expand the label hierarchy in Fig. 2(a) by adding other attribute nodes (e.g., race and attractiveness). Additionally, the FFSC dataset could integrate other modalities, such as text, to enable multimodal face forgery detection. In such cases, multimodal large-language models could be used to generate semantically localized descriptions of face manipulations [89].
@@
Our current definition of face forgery is limited to face manipulation methods. It would be beneficial to expand this definition to include face creation methods, such as those utilizing GANs [3], diffusion models [2, 91, 92] and computer graphics [93]. However, this expansion presents significant challenges. First, technology has advanced to the point where we can render an existing face presented in a real physical scene with such realism that it becomes indistinguishable from an actual photograph (see Fig. 8). Second, there is evidence that generative models can memorize training data [94], and techniques exist to retrieve these memorized images (see Fig. 9). In both cases, determining the authenticity of a face image requires heightened awareness or insights to perceive its true nature. Taking a step back, a more pragmatic extension of our definition is to include digital manipulation methods for images of natural scenes that do not necessarily contain faces.
Figure 9: Memorization results [94] of a diffusion model trained on a small-sized CelebA dataset [95]. The top row depicts training face images downsampled to $80\times 80$, and the bottom row displays face images generated by the diffusion model, which closely resemble those from the training set.
Acknowledgements
The authors would like to thank Mr. Weiran Zhao for his kind help in data collection. This work was supported in part by the Hong Kong RGC General Research Fund (11220224), the CityU Strategic Research Grants (7005848 and 7005983), and an Industry Gift Fund (9229179).